OpenStudy (anonymous):

Solve S = 4v2 for v show work

OpenStudy (anonymous):

\[S = 4v^2 \rightarrow \frac{S}{4} = v^2 \rightarrow v= \sqrt {\frac{S}{4}} \] \[v= {\frac{\sqrt S}{2}}\]